MODS1 checkout (2017 October 30, first half)

MODS1 was installed on Saturday, 2017 October 21.

Two issues arose in starting up the instrument during the week and these were resolved:
  • The HEB_B CSR needed to be replaced IT 6866
  • The Red Grating microLynx controller failed and the cable was swapped to use a spare microLynx IT 6877
Another issue was noted --- a dirty field lens --- and this remains to be resolved.

IMCSTTF zeropoints were updated on Saturday and checked at the beginning of the night.

ISA: Olga

SO: Jenny

TO: David

Summary on MODS checkout:

  • delta theta is close to what it historically has been, ~-0.2 deg, which suggests that the rotator zeropoint does not need to be changed. We did MOS acquisitions on 2 masks: StoneO (red and blue acq) and StoneA (red acq). We can use the old value in PCSInstrument.conf and the old transform for MODS1.

  • Stars are seen in boxes in the first thru-mask image, so pointing origin in PCSInstrument.conf file is OK.

  • BD+284211 observed in all modes. Feige110 observed in dual prism/grating and red prism/grating. This tested 5\x94 slit acquisition and got useful standard star data on a clear (and I think photometric) night.

  • Dice5 repeated 3 times on Stone O (though images looked elongated earlier, these look OK)

  • Long-slit alignment to be done

  • MODS1 looks OK (but has cataracts\x85 the debris on the field lens is really ugly. We confirm that the Y=620 position for MODS1 long-slit spectroscopy puts the star well below the gunk. I hope it can be cleaned.

4:30 MODS1 mask snap, followed by imaging flats. Only got u before quit so we could do test preset. U_sdss mods1b.20171030.0008-12

5:40 Running IMCS Lock test/confirmation imcsTTFs.pro for MODS1 @60deg rot 207. (This was run yesterday evening and b/rconfig.tab tables were updated with new values which were similar to the old except for red dual prism, where yesterday imcslock timed out on first try.).

IMCSlock achieved on first try for all modes. Mods1b.20171030.00013-18 mods1r.20171030.0008-13

5:56 Opening

6:05 Pointing and collimation

6:24 rm507-3 froze up during StoneO target acquisition. Olga was able to take over NX session to obs2 on rm507-5.

6:20 Stone O rotator center PA=0 active script send for collimation. Always doing instconfig to lock IMCS but guiding only at first PA.

Active PA0: mods1b 19 mods1r 14 (elongated stars on-axis; wfs shows some evidence of spherical)

Track PA180: mods1b 20 mods1r 15

Rotcen from 19&20 -> x0,y0= (1555.6,1565)

Rotcen from 15&16 -> x0,y0= (1550,1570)

Track PA90: mods1b 21 mods1r 16

Track PA270: mods1b 22 mods1r 17

21&22 -> x0,y0 = (1539,1541)

16&17 -> x0,y0 = (1536,1547) unsure why the 90/270 values do not agree with 0/180 ones, but I saw this before\x85 offline will have another look. (Dave solved this! Both observations need to be track mode presets. So the true rotator center are the ones above for 21&22 and 16&17).

6:47 resending active preset at 90 to recollimate

Active PA90: mods1b 23 mods1r 18

Track PA270: mods1b 24 mods1r 19

18&19 -> x0,y0 = (1569,1537) but the geomap residuals are ~1-2 pix in x and 0.5 pix in y\x85

preset posangle B R
active 0 19 14
track 180 20 15
track 90 21 16
track 270 22 17
active 90 23 18
track 270 24 19

Blue Rotator Centers
pair dtheta xrc yrc Comments
20 & 21 -90 1533.3 1538.6
20 & 22 +90 1541.8 1535.1
21 & 22 +180 1539.5 1541.0
21 & 20 +90 1533.4 1538.5
22 & 20 -90 1541.8 1535.1
22 & 21 -180 1539.4 1541.0
spread all 8.5 pix 5.9 pix
    1.02" 0.708"

Red Rotator Centers
pair dtheta xrc yrc Comments
15 & 16 -90 1528.9 1544.5
15 & 17 +90 1539.3 1539.3
16 & 17 +180 1537.0 1547.1
Reverse order yields same results, to within 0.1-0.2 pixels
spread all 10.1 pix 7.8 pix
    1.242" 0.959"

07:05 MST: Stone O MOS acquisition

R 20, 21

modsAlign - used all boxes and stars except for the two under the guide probe:

Offsetpointing -0.1867 -0.776 1.361 detxy rel

Fixed problem with mask table (updateSlits needed to be run to update mask table on MODS1 IE server)

New acquisition on Stone O field: 23 & 24

modsAlign now with only selected boxes, to avoid the two vignetted by the guide probe: 23 & 24

Offsetpointing -0.1815 -0.940 1.074 detxy rel

Confirmatory image looks pretty good 25

All stars appear elongated --- is this thermal? Spherical? David is calling AO support.

Meanwhile, we are trying a blue acquisition on this field.

19:45 Blue acq only Acq images 27, 28 mods1b.

Offsetpointing -0.2153 -0.951 0.973 detxy rel

29 confirmatory thru-slit image, looks good.

19:56 Sending dice 5 PA=0 script on stoneO

Detxy absolute offsets 20\x94
DetXY Abs x4: (-20,20),(20,-20),(-20,-20),(20,20),(0,0) mods1b 0030-44 & mods1r 0026-40

COMM error (setting \x93partner\x94 timed out)

20:30 Going to BD+28. Nice and clear. Taking standard in all modes. Airmass 1.023 and setting modsAlign offset (-0.417, 10.616). Reran modsAlign on confirmation image 43 and offset (-0.465, -0.070), Seeing just jumped up. 1.35 on guider and 1.8 on DIMM.

BD28 Acquisition Mods1r 41-43

BD28 Dual Grating mods1b 45-47 mods1r 44-46

BD28 Red Grating mods1r 47-49

BD28 Blue Grating mods1b 48-50

BD28 Dual Prism mods1b 51-53 mods1r 50-52

BD28 Red Prism mods1r 53-55

BD28 Blue Prism mods1b 54-56

21:10 bouncing between 0.8 and 1.85\x94 on guider, highly variable.

21:29 StoneA mos acquisition to confirm rotator zeropoint, mosAlign. 056-057 slit and sky image. Offset -0.1897, -0.671 0.333 detxy rel

21:41 Sent preset for Feige110. Seeing ~1.25\x94, but varying on short timescales from 0.9 to 1.8



Offsets -1.139, 11.189. Reran confirmation modsAlign. Offsets 0.088, -0.022

Feige110 Acquisition Mods1r 59-60

Feige110 Dual Grating mods1b mods1r 61-62

Feige110 Red Grating mods1r 63-64

Feige110 Blue Grating mods1b

Feige110 Dual Prism mods1b 59-60 mods1r 67-68

Feige110 Red Prism mods1r 65-66

22:14 David is doing a pointing check

22:18 Sent preset to Al\x92s NSI target. Forgot to change to pseudo mono mode. David is doing that now for us.

22:33 Mods acquisition sent. Failed because no guide star in patrol field

22:44 LBC data being taken. Going to try to manually take exposures for mods

22:48 Cannot manually take spectra because we cannot send offsets without a preset. LBC data is more important at the moment. Unable to guide non-sidereally on this target with mods. Guide star just flies out of the FOV

22:55 can

23:07 Decided to just manually take unguided images 300sec in g and i mods1b 61-62, mods1r 72-73. LBC data being taken in z simultaneously guided.

23:25 Right Primary nearly in limit. Decided to change from pseudo monocular to monocular lbc, clear active optics.

23:47 Took garbage slave mods images. MODS was not tracking, rotator not rotating. Mods1b 63-64 and mods1r 74-75.

23:51 Preparing to handover telescope. LBC images continue

0:10 Done taking LBC exposures, Emanuele is all set up for MODS1. He had authorization issues with LBTB and mods (ssh keys for LBTB account), so Olga set him up on the LBTO account

0:12 David has completed the reauthorization


Both VFLAT lamps on MODS1 and MODS2 were changed last summer. Calibrations with MODS1 gave the following count levels (as summarized by JPower):

mods1r.20171026.000[1-7].fits object maskname gratname callamps filtname vflat exptime max counts
mods1r.20171026.0002.fits VFLAT + r_sdss Dual Imaging Flat Imaging Flat VFLAT r_sdss 3.0 2.0 25K
mods1r.20171026.0003.fits VFLAT + i_sdss Dual Imaging Flat Imaging Flat VFLAT i_sdss 1.5 7.0 20K
mods1r.20171026.0004.fits VFLAT + z_sdss Dual Imaging Flat Imaging Flat VFLAT z_sdss 1.5 3.0 20K
mods1r.20171026.0005.fits QTH1 + ND1.5 + ND1.5 Spectral Flat LS5x60x1.0 G670L QTH6V QTH1 ND1.5 1.5 1.0 27K
mods1r.20171026.0006.fits VFLAT + Clear Spectral Flat LS5x60x1.0 G670L VFLAT Clear 10.0 1.5 23K
mods1r.20171026.0007.fits VFLAT + Clear + ND1.5 Spectral Flat Imaging G670L VFLAT ND1.5 5.0 1.5 33K
 
mods1b.20171026.0002.fits VFLAT + u_sdss Dual Imaging Flat Imaging Flat VFLAT u_sdss 10.0 50.0 23K
mods1b.20171026.0003.fits VFLAT + g_sdss Dual Imaging Flat Imaging Flat VFLAT g_sdss 4.0 12.0 28 K
mods1b.20171026.0004.fits QTH1 + ND1.5 + ND1.5 Spectral Flat LS5x60x1.0 G400L QTH6V QTH1 ND1.5 1.5 16.0 31K
mods1b.20171026.0005.fits QTH1 + UG5 Spectral Flat LS5x60x1.0 G400L QTH6V QTH1 UG5 1.5 7.0 27K
mods1b.20171026.0006.fits VFLAT + Clear + ND1.5 Spectral Flat Imaging G400L VFLAT Clear 5.0 5.0 37K
mods1b.20171026.0007.fits VFLAT + UG5 Spectral Flat Imaging G400L VFLAT UG5 10.0 2.0 32K
Series of biases at 8Kx3K, 3Kx3K, 4Kx3K and 1Kx1K followed (these were both MODS1 and MODS2).

Need to edit slitflats_r 1.2" slitflat to use exptime 1.5-sec, not 1.0-sec which gives 18K.

Observer ran red-only calibrations:
  • grpixflats_r.cal,
  • slitflats(1.2")_r.cal and
  • slitflats_5as_r.cal.

I later ran:
  • grpixflats.cal (dual):
    • red Vflat 5 t=1.5s gives 38K counts (132) consider setting t=1 instead
    • blue Vflat 10 UG5 t=2 32K (72)
  • grlamps (red, dual, blue)
  • grpixflats_b (blue only)

- %USERSIG{OlgaKuhn - 2017-10-30}%
Topic revision: r4 - 30 Nov 2017, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback